OpenStudy (anonymous):

Estimate to the nearest integer. squareroot 14

Parth (parthkohli):

Square root of 14 is between \(\sqrt9\) i.e 3 and \(\sqrt {16}\) i.e 4. It looks much more nearer to \(\sqrt{16}\)

Parth (parthkohli):

What is \(\sqrt{16}\) in integer form?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

squareroot 27 there is that one to it says Estimate to the nearest integer.

Parth (parthkohli):

It looks nearer to \(\sqrt{25}\). What is the square root of 25?

OpenStudy (anonymous):

um i don't know i am barley learning this stuff....